# Infusion: internal diffusion for inpainting of dynamic textures and complex motion

### v3

@@
Abstract.
Video inpainting is the task of filling a region in a video in a visually convincing manner. It is very challenging due to the high dimensionality of the data and the temporal consistency required for obtaining convincing results. Recently, diffusion models have shown impressive results in modeling complex data distributions, including images and videos. Such models remain nonetheless very expensive to train and to perform inference with, which strongly reduce their applicability to videos, and yields unreasonable computational loads. We show that in the case of video inpainting, thanks to the highly auto-similar nature of videos, the training data of a diffusion model can be restricted to the input video and still produce very satisfying results. This leads us to adopt an internal learning approach, which also allows us to greatly reduce the neural network size by about three orders of magnitude less than current diffusion models used for image inpainting. We also introduce a new method for efficient training and inference of diffusion models in the context of internal learning, by splitting the diffusion process into different learning intervals corresponding to different noise levels of the diffusion process. To the best of our knowledge, this is the first video inpainting method based purely on diffusion. Other methods require additional components such as optical flow estimation, which limits their performance in the case of dynamic textures and complex motions. We show qualitative and quantitative results, demonstrating that our method reaches state of the art performance in the case of dynamic textures and complex dynamic backgrounds.
inpainting, video, diffusion models, editing
submissionid: papers_1080ccs: Computing methodologies Image processing
Figure 1. Our method can inpaint videos with dynamic textures and complex motion using a diffusion-based approach (inpainted region in blue in the film reel). High quality results are obtained with a reasonable network size by learning on the input video only, and leveraging our proposed training strategy: interval training. Training and inference are mixed during the diffusion process, which is divided into intervals.
@@ Abstract.
A video inpainting example of our method, with a person in front of a fireplace inpainted. Close-up view of the fire reveals that the completion is realistic and dynamic.
1. Introduction
Video inpainting, or video completion, aims to fill in an unknown region in a video, in a visually convincing manner. While in the case of images, very impressive results have been obtained using deep learning approaches (Yu et al., 2018; Saharia et al., 2022a), the extension of these methods to video has proven to be extremely difficult. This is due to the great technical challenges of videos, such as the large dimensionality of the data which leads to computationally expensive methods, the high sensitivity of the human eye to temporal inconsistencies, and the difficulty of inpainting complex motions. Many current state of the art methods combine deep learning with optical flow to deal with temporal consistency and motion. Unfortunately, optical flow breaks down in the presence of complex motions, for example in the case of dynamic (video) textures, which are poorly synthesized by such methods, as we show in this paper.
Concurrently, diffusion models have recently emerged as a powerful framework to produce samples of complex distributions, allowing the synthesis of realistic and diverse examples of images, with unprecedented quality (Ho et al., 2020). However, these networks are extremely large (up to 500 million parameters for image inpainting (Lugmayr et al., 2022; Saharia et al., 2022a)), leading to very long training and inference times. While several acceleration techniques have been proposed (Song et al., 2021; Rombach et al., 2022) in the case of images, they would nevertheless yield prohibitive training times for video inpainting.
In this paper, we propose a diffusion-based video inpainting method which has both a reasonable network size (around 500k parameters), and avoids relying on optical flow that can lead to poor reconstruction of complex motions. We show that a lightweight diffusion network, combined with an “internal learning” method, produces high-quality video inpainting results, especially in the difficult case of complex motion. Internal learning refers to restricting the training database to the input video, which makes sense, due to the highly auto-similar nature of videos (Wexler et al., 2007). In particular, we propose a novel training algorithm for diffusion models, referred to as “interval training”, which allows us to greatly reduce the size of the network, leading to feasible training and inference times (between 3-15 hours on a single GPU). Furthermore, to the best of our knowledge, our work yields the first purely diffusion-based video inpainting algorithm, we avoid using the aforementioned optical flow. Thus, in contrast to most of the other video inpainting algorithms, our method handles dynamic textures well and is able to produce diverse inpainting solutions. Finally, temporal coherence is ensured by 3D convolutions in the network.
@@ 1. Introduction
2. Related work
2.1. Diffusion Models
@@
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have shown great generative capabilities on image datasets. They have supplanted Generative Adversarial Networks (GAN) in quality and training stability (Dhariwal and Nichol, 2021). One major drawback, however, is the slow inference and training, which has lead to several attempts at acceleration up with various techniques (Rombach et al., 2022; Nichol and Dhariwal, 2021).
For image inpainting, Lugmayr et al. (2022) use an unconditional diffusion model, with projections and time-travels for better coherence. Saharia et al. (2022a) train a conditional model for diverse image-to-image problems.
However, the extension of diffusion models to videos is not straightforward, in particular due to their huge sizes and training times. Blattmann et al. (2023) finetune an image model for videos by introducing temporal attention layers. Another option is to decompose the video as multiple images via neural atlases (Kasten et al., 2021) which can then be handled by image-based diffusion models (Chai et al., 2023). Ceylan et al. (2023) use noise inversion and attention maps for consistent video editing.
For video generation, Ho et al. (2022b) have extended the image diffusion approach by using a temporal attention layer for sharing the information along the temporal axis, avoiding costly 3D convolutions. Other approaches have studied the conditional generation of videos via multiple networks for frame interpolation and super-resolution (Ho et al., 2022a; Singer et al., 2023). Harvey et al. (2022) implement a diffusion network for generation based on long temporal attention and frame interpolation. To reduce the computational complexity of handling full 3D videos, Yu et al. (2023) use multiple 2D projections. Mei and Patel (2023) have two models for frame content and frame motion. The very recent diffusion model Sora produces videos from text with unprecedented quality. While the complete architecture and training strategy is not known at the time of writing (Brooks et al., 2024), it is likely that training or even running such a model requires unreasonable computing facilities.
@@
Balaji et al. (2023) propose to create an ensemble of diffusion networks specialised in different steps of the diffusion process. This approach shares some similarities with our training method (it was developed concurrently to our work), however a major difference is that theirs does not address the issue of network size, which is so crucial in our video inpainting problem.
@@
Related to the internal learning strategy proposed in our work, it has also been shown recently that diffusion can be applied in a small data setting, for instance when only a single image is available for training. Kulikov et al. (2023) thus achieve single image generation using only one image. Similarly, Nikankin et al. (2023) show the application to image as well as video (single video generation / extrapolation). It is also possible to finetune a general diffusion model for a specific scene (Ruiz et al., 2023).
2.2. Video Inpainting
@@
Classical approaches to video inpainting have long relied on available information already present in the video to achieve inpainting of the occluded region. They rely on the auto-similarity hypothesis, which is the foundation for patch-based methods in inpainting (Wexler et al., 2007; Newson et al., 2014), as well as other problems such as denoising (Buades et al., 2005) or image editing (Pritch et al., 2009).
@@
Another class of approaches is based on optical flow. Indeed, in many situations, especially when the occluded region is observed at some point in time in the video, optical flow gives very useful information. Huang et al. (2016) jointly optimize the appearance and the optical flow, making sure that the output is temporally coherent.
@@
More recently, Kim et al. (2019) have proposed a direct inpainting approach using a 3D neural network. Methods such as (Xu et al., 2019; Gao et al., 2020) first inpaint the optical flow and then propagate pixel values. These methods have been successful in handling non-rigid motions and long range dependencies. Zhang et al. (2022c) proposed to predict the optical flow using more frames, resulting in a more accurate optical flow and better preservation of motions. Li et al. (2022) integrate the optical flow prediction and propagation inside the training loop, optimizing the network parameters for video inpainting end-to-end.
@@
Transformer-based architectures have also been used for video inpainting (Liu et al., 2021; Zhang et al., 2022a). Transformers can handle long range dependencies needed in video inpainting, however, their heavy computational cost can be prohibitive for the large dimensionality of videos and imply drastic approximations of the attention. ProPainter (Zhou et al., 2023a) improves the long range propagation of known regions with Transformers.
@@
Internal learning for video inpainting has been proposed by Zhang et al. (2019). They learn a convolutional model on a single video similarly to Deep Image Prior (Ulyanov et al., 2018), and model the optical flow for better temporal coherency. Ouyang et al. (2021) tackle the same problem but propose a solution to avoid explicit modeling of the optical flow. Finally, Ren et al. (2022) use a Transformer with a discrete codebook finetuned on the target video, similar to a masked quantized autoencoder.
Figure 2. Our UNet architecture used for denoising the video in our diffusion approach is very simple, using mainly 3D convolutions.
[…]
Our UNet architecture uses 3D convolutions, downsampling and upsampling operations. The number of layers is low and is easy to illustrate.
3. Method
3.1. Background
@@
Figure 3. We train a diffusion model on the input video only using interval training. In interval training, we learn on a subset of timesteps and the inference is done immediately after the training phase. Starting from $x_{T}$, the inpainting result is progressively generated with a perfectly adapted network.
@@
We first recall the diffusion framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), based on discrete forward and backward processes. We adapt it for inpainting by modeling the conditional data distribution $p(\mathrm{x}\mid y)$ where $\mathrm{x}$ is a full video, and $y$ are the observations i.e. the known (unoccluded) region of the video. We represent an RGB video of $L$ frames of size $H\times W$ with a vector in $\mathbb{R}^{L\times H\times W\times 3}$. Starting from the unknown distribution of completed videos $q(\mathrm{x}_{0}\mid y)$ conditioned on $y$, we define a Markov chain $q(\mathrm{x}_{0}\mid y),...,q(\mathrm{x}_{T}\mid y)$ with the forward transition kernel for $t\in[1,T]$:
(1) $q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1},y)=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\bm{I}\right),$
@@
with the variance schedule linearly increasing from $\beta_{1}=0.0001$ to $\beta_{T}=0.02$. Associated with the forward process is the reverse process defining the data distribution $p_{\theta}(\mathrm{x}_{t}\mid\mathrm{x}_{t+1},y)$ which is also a Gaussian of unknown but learnable mean and variance:
@@
We the classical approach of a neural network to predict the conditional mean $\mu_{\theta}(\mathrm{x}_{t+1},y,t)$ whose parameters $\theta$ are optimized to maximize a lower bound of the log-likelihood of the data. Following (Ho et al., 2020) we use a fixed variance schedule for the reverse process: $\sigma^{2}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$.
In the case of inpainting, we are interested in the conditional data distribution $p(x\mid y)$ where $y$ are our observations i.e. the known region of the video, we thus train our network with additional inputs. We minimize the L2 loss of the $x_{0}$-parameterization (Ho et al., 2020):
(2) $\mathbb{E}_{\mathrm{x}_{0},\epsilon,t,M}\lVert M\circ\left(\mathrm{x}_{0}-f_{\theta}(\tilde{\mathrm{x}}_{t},y,t,M)\right)\rVert_{2}^{2},$
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\Pi_{i=0}^{t}\alpha_{t}$, $\epsilon\sim\mathcal{N}(0,I)$. The observations and masks are introduced via concatenation in the input layer (similarly to (Saharia et al., 2022a)).
Unless specified, we use the default parameters of (Ho et al., 2020): variance schedule, sampling, time schedule. Therefore, our method could benefit from many of the recent improvements for each of these items.
3.2. Architecture
For temporal coherency, we extend the traditional UNet architecture to videos by using 3D convolutions. Another option is to rely on optical flow, but this can be inaccurate for complex situations. Another common approach for long-range dependencies is to use the attention mechanism (Vaswani et al., 2017), but it is extremely computationally heavy, and prohibitively so in the case of videos. Our methods avoids the use of this mechanism, yielding a network with reasonable computational requirements.
Since we are only training the model on a single video, we can greatly reduce the number of parameters in our network in comparison to standard diffusion models. The network has a total of 16 convolutional layers, distributed over 4 levels. Our 3D UNet downsamples only in the spatial dimensions to preserve as much as possible the temporal dimension, which is already much smaller than the others. We only use 32 output channels in each layer resulting in only 500k parameters which are to be compared to the 500M parameters of Palette (Saharia et al., 2022a) for image inpainting and the billions of parameters of other video diffusion models (Ho et al., 2022a). Our code can be found in the supplementary materials.
3.3. Single Video (Internal) Training
Our neural network is trained to inpaint a single video, using the video itself. This type of approach is referred to as internal learning. Fundamentally, it is based on the auto-similarity hypothesis, which has been successfully used in the case of patch-based methods (Newson et al., 2014) and neural networks via internal learning (Zhang et al., 2019; Ouyang et al., 2021). The main assumption here, explaining the success of such methods for video inpainting, is that videos contain roughly the same content in successive frames, with some motion. with some or all of the occluded content being revealed at some point in time. Thus, what might be considered as overfitting in a normal machine learning situation is in fact desirable here: we want to use the content of the video to inpaint itself. Put another way, we take advantage of both the learning and auto-similarity paradigms in our method. This kind of situation also arises in the problem of Single-Image Generation  (Shaham et al., 2019).

### v4

@@
Abstract
Video inpainting is the task of filling a region in a video in a visually convincing manner. It is very challenging due to the high dimensionality of the data and the temporal consistency required for obtaining convincing results. Recently, diffusion models have shown impressive results in modeling complex data distributions, including images and videos. Such models remain nonetheless very expensive to train and to perform inference with, which strongly reduce their applicability to videos, and yields unreasonable computational loads. We show that in the case of video inpainting, thanks to the highly auto-similar nature of videos, the training data of a diffusion model can be restricted to the input video and still produce very satisfying results. With this internal learning approach, where the training data is limited to a single video, our lightweight models perform very well with only half a million parameters, in contrast to the very large networks with billions of parameters typically found in the literature. We also introduce a new method for efficient training and inference of diffusion models in the context of internal learning, by splitting the diffusion process into different learning intervals corresponding to different noise levels of the diffusion process. We show qualitative and quantitative results, demonstrating that our method reaches or exceeds state of the art performance in the case of dynamic textures and complex dynamic backgrounds.
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010382.10010383</concept_id> <concept_desc>Computing methodologies Image processing</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@ Abstract
[…]
volume: 44issue: 2
1 Introduction
Video inpainting, or video completion, aims to fill in an unknown region in a video, in a visually convincing manner. While in the case of images, very impressive results have been obtained using deep learning approaches [YLY18, SCC22], the extension of these methods to video has proven to be extremely difficult. This is due to the great technical challenges of videos, such as the large dimensionality of the data which leads to computationally expensive methods, the high sensitivity of the human eye to temporal inconsistencies, and the difficulty of inpainting complex motions. Many current state of the art methods combine deep learning with optical flow to deal with temporal consistency and motion [LLQ22, ZLCC23]. Unfortunately, optical flow breaks down in the presence of complex motions, for example in the case of dynamic (video) textures such as waves or tree leaves, which are poorly synthesized by such methods, as we show in this paper.
Concurrently, diffusion models have recently emerged as a powerful framework to produce samples of complex distributions, allowing the synthesis of realistic and diverse examples of images, with unprecedented quality [HJA20]. However, these networks are extremely large (up to 500 million parameters for image inpainting [LDR22, SCC22]), leading to very long training and inference times. While several acceleration techniques have been proposed [SME21, RBL22] in the case of images, they would nevertheless yield prohibitive training times for video inpainting. Furthermore, while the common practice of fine-tuning large diffusion models for specific tasks is feasible in the case of images, it is still not practical in the case of videos due to the large size of video diffusion models.
In this paper, we propose a diffusion-based video inpainting method which has both a reasonable network size (around 500k parameters), and avoids relying on optical flow that can lead to poor reconstruction of complex motions. We show that a lightweight diffusion network, combined with an “internal learning” method, produces high-quality video inpainting results, especially in the difficult case of complex motion. Internal learning refers to restricting the training database to the input video, which makes sense, due to the highly auto-similar nature of videos [WSI07]. In particular, we propose a novel training algorithm for diffusion models, referred to as “interval training”, which significantly improves the performance in this context. This allows us to generate high quality results with a lightweight network, leading to feasible training and inference times (between 3-15 hours on a single GPU). In contrast to most of the other video inpainting algorithms which mainly rely on optical flow, our method handles dynamic textures well and is able to produce diverse inpainting solutions. Finally, temporal coherence is ensured by 3D convolutions in the network.
@@ 1 Introduction
2 Related work
2.1 Diffusion Models
@@
Diffusion models [SWMG15, HJA20] have shown great generative capabilities on image datasets. They have supplanted Generative Adversarial Networks (GAN) in quality and training stability [DN21]. One major drawback, however, is the slow inference and training, which has lead to several attempts at acceleration up with various techniques [RBL22, ND21].
@@
Balaji et al. [BNH23] propose to create an ensemble of diffusion networks specialised in different steps of the diffusion process. This approach shares some similarities with our training method (it was developed concurrently to our work), however a major difference is that theirs does not address the issue of network size, which is so crucial in our video inpainting problem.
For image inpainting, Lugmayr et al. [LDR22] use an unconditional diffusion model, with projections and time-travels for better coherence. Saharia et al. [SCC22] train a conditional model for diverse image-to-image problems. User-guided inpainting is possible with text [XZL22] or other modalities (strokes, examples) [YCL23].
However, the extension of diffusion models to videos is not straightforward, in particular due to their huge sizes and training times. Blattmann et al. [BRL23] finetune an image model for videos by introducing temporal attention layers. Another option is to decompose the video as multiple images via neural atlases [KOWD21] which can then be handled by image-based diffusion models [CGWL23]. Ceylan et al. [CHM23] use noise inversion and attention maps for consistent video editing.
For video generation, Ho et al. [HSG22] have extended the image diffusion approach by using a temporal attention layer for sharing the information along the temporal axis, avoiding costly 3D convolutions. Other approaches have studied the conditional generation of videos via multiple networks for frame interpolation and super-resolution [HCS22, SPH23]. Harvey et al. [HNM22] implement a diffusion network for generation based on long temporal attention and frame interpolation. To reduce the computational complexity of handling full 3D videos, multiple 2d projections can be used [YSKS23]. Mei et al. [MP23] have two models: one for frame content and one for frame motion. The very recent diffusion model Sora [BPH24] produces videos from text with unprecedented quality. While the complete architecture and training strategy is not known at the time of writing , it is likely that training or even running such a model requires unreasonable computing facilities.
@@
Related to the internal learning strategy proposed in our work, it has also been shown recently that diffusion can be applied in a small data setting, for instance when only a single image is available for training. Kulikov et al. [KYKM23] thus achieve single image generation using only one image. Similarly, Nikankin et al. [NHI23] show the application to image as well as video (single video generation / extrapolation). It is also possible to finetune a general diffusion model for a specific scene [RLJ23].
2.2 Video Inpainting
@@
Classical approaches to video inpainting have long relied on available information already present in the video to achieve inpainting of the occluded region. They rely on the auto-similarity hypothesis, which is the foundation for patch-based methods in inpainting [WSI07, NAF14], as well as other problems such as denoising [BCM05] or image editing [PKP09].
@@
Another class of approaches is based on optical flow. Indeed, in many situations, especially when the occluded region is observed at some point in time in the video, optical flow gives very useful information. [HKAK16] jointly optimize the appearance and the optical flow, making sure that the output is temporally coherent.
@@
More recently, [KWLK19] have proposed a direct inpainting approach using a 3D neural network. Methods such as [XLZL19, GSHK20] first inpaint the optical flow and then propagate pixel values. These methods have been successful in handling non-rigid motions and long range dependencies. [ZFL22c] proposed to predict the optical flow using more frames, resulting in a more accurate optical flow and better preservation of motions. [LLQ22] integrate the optical flow prediction and propagation inside the training loop, optimizing the network parameters for video inpainting end-to-end.
@@
Transformer-based architectures have also been used for video inpainting [LDH21, ZFL22a]. Transformers can handle long range dependencies needed in video inpainting, however, their heavy computational cost can be prohibitive for the large dimensionality of videos and imply drastic approximations of the attention. ProPainter [ZLCC23] improves the long range propagation of known regions with Transformers.
@@
Internal learning for video inpainting has been proposed by [ZMJ19]. They learn a convolutional model on a single video similarly to Deep Image Prior [UVL18], and model the optical flow for better temporal coherency. [OWC21] tackle the same problem but propose a solution to avoid explicit modeling of the optical flow. Finally, [RZZ22] use a Transformer with a discrete codebook finetuned on the target video, similar to a masked quantized autoencoder.
Only a few works tackle video inpainting using diffusion models, due to their computational requirements. AVID [ZWW24], and CoCoCo [ZZQ24] propose different solutions to text-guided video inpainting, which is a different problem than our own i.e. automatic video inpainting without user intervention or guidance. [LCS24] combine a diffusion model with optical flow for temporal consistency, however the method is still limited when the optical flow is inaccurate.
3 Method
3.1 Background
@@
Figure 1: We train a diffusion model on the input video only using interval training. In interval training, we learn on a subset of timesteps and the inference is done immediately after the training phase. Starting from $x_{T}$, the inpainting result is progressively generated with a perfectly adapted network.
@@
We first recall the diffusion framework of Denoising Diffusion Probabilistic Models (DDPM) [HJA20], based on discrete forward and backward processes. We adapt it for inpainting by modeling the conditional data distribution $p(\mathrm{x}\mid y)$ where $\mathrm{x}$ is a full video, and $y$ are the observations i.e. the known (unoccluded) region of the video. We represent an RGB video of $L$ frames of size $H\times W$ with a vector in $\mathbb{R}^{L\times H\times W\times 3}$. Starting from the unknown distribution of completed videos $q(\mathrm{x}_{0}\mid y)$ conditioned on $y$, we define a Markov chain $q(\mathrm{x}_{0}\mid y),...,q(\mathrm{x}_{T}\mid y)$ with the forward transition kernel for $t\in[1,T]$:
$q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1},y)=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\bm{I}\right),$ (1)
@@
and the variance schedule linearly increasing from $\beta_{1}=0.0001$ to $\beta_{T}=0.02$. Associated with the forward process is the reverse process defining the data distribution $p_{\theta}(\mathrm{x}_{t}\mid\mathrm{x}_{t+1},y)$ which is also a Gaussian of unknown but learnable mean and variance:
@@
We use the classical approach of a neural network to predict the conditional mean $\mu_{\theta}(\mathrm{x}_{t+1},y,t)$ whose parameters $\theta$ are optimized to maximize a lower bound of the log-likelihood of the data. Following [HJA20] we use a fixed variance schedule for the reverse process: $\sigma^{2}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$.
In the case of inpainting, we are interested in the conditional data distribution $p(x\mid y)$ where $y$ are our observations i.e. the known region of the video, we thus train our network with additional inputs. We minimize the L2 loss of the $x_{0}$-parameterization [HJA20]:
$\mathbb{E}_{\mathrm{x}_{0},\epsilon,t,M}\lVert M\circ\left(\mathrm{x}_{0}-f_{\theta}(\tilde{\mathrm{x}}_{t},y,t,M)\right)\rVert_{2}^{2},$ (2)
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\Pi_{i=0}^{t}\alpha_{t}$, $\epsilon\sim\mathcal{N}(0,I)$. The observations and masks are introduced via concatenation in the input layer (similarly to [SCC22]).
Unless specified, we use the default parameters of [HJA20]: variance schedule, sampling, time schedule. Therefore, our method could benefit from many of the recent improvements for each of these items.
3.2 Architecture
To ensure temporal coherency, we extend the traditional UNet architecture to videos by using 3D convolutions. Another option is to rely on optical flow, but this can be inaccurate for complex situations such as dynamic textures or complex motion with severe occlusions. Another common approach for long-range dependencies is to use the attention mechanism [VSP17], but it is extremely computationally heavy, and prohibitively so in the case of videos. Our methods avoids the use of this mechanism, yielding a network with reasonable computational requirements.
Since we are only training the model on a single video, we can greatly reduce the number of parameters in our network in comparison to standard diffusion models. The network has a total of 16 convolutional layers, distributed over 4 levels. All convolutions are 3x3x3, with padding 1, stride 1. Downsampling is achieved using a max pooling operation while upsampling uses nearest neighbor interpolation. Feature maps from skip connections and feature maps from upsampling layers are combined using concatenation. Our 3D UNet downsamples only in the spatial dimensions to preserve as much as possible the temporal dimension, which is already much smaller than the others. The full architecture details are available in Appendix, Table 7. We only use 32 output channels for all internal layers resulting in only 500k parameters which are to be compared to the 500M parameters of Palette [SCC22] for image inpainting and the billions of parameters of other video diffusion models [HCS22]. Our code can be found in the supplementary materials and is available at https://github.com/ncherel/infusion.
3.3 Single Video (Internal) Training
Our neural network is trained to inpaint a single video, using the video itself. This type of approach is referred to as internal learning. Fundamentally, it is based on the auto-similarity hypothesis, which has been successfully used in the case of patch-based methods [NAF14] and neural networks via internal learning [ZMJ19, OWC21]. The main assumption here, explaining the success of such methods for video inpainting, is that videos contain roughly the same content in successive frames, with some motion. with some or all of the occluded content being revealed at some point in time. Thus, what might be considered as overfitting in a normal machine learning situation is in fact desirable here: we want to use the content of the video to inpaint itself. Put another way, we take advantage of both the learning and auto-similarity paradigms in our method. This kind of situation also arises in the problem of Single-Image Generation  [SDM19].